# Distributed memory Tensor-train crossT. Shi, D. Hayes, and J. Qiu

### v1

@@ 3.1 Submatrix parallel matrix cross approximation
3.1.1 A new iterative construction of matrix cross approximation
@@
where $\tilde{E}_{z}:=\tilde{A}_{z}-A$. A similar form of Eq. 7 can be found in [cortinovis2020maximum], where the construction is led by an LU decomposition of the approximation. In other words, the construction starts with full data of the original matrix. On the contrary, ours begins with a zero initialization of the approximation.
As an example, we perform a greedy cross approximation on a 100$\times$100 Hilbert matrix defined elementwise via
@@
which is shown to have rapidly decaying singular values in [beckermann2019bounds]. The behaviors of Eq. 7 and the direct computation for approximation construction is shown in Fig. 2. There we can see that the use of formula Eq. 7 provides better numerical stability up to machine precision, while the direct method suffers from degeneracy when selected rank is larger than 12. A separate advantage of Eq. 7 is that we are able to make global approximation updates at a local level, which is essential for communication efficiency of submatrix and subtensor algorithms discussed in the following subsections.
Figure 2: Relative Error versus rank using formula (7) and a direct computation of the factor matrices. The test was run a 100$\times$100 Hilbert matrix.
@@ 3.1.2 Iterative construction of submatrix cross approximation
2:A new pivot $(i^{*},j^{*})$ for the entire matrix $A$.
@@ 3.1.3 An alternative iterative construction of matrix and submatrix cross approximation
For submatrix ACA, Eq. 9 is sufficient for both finding pivots and constructing approximations. However, building TT cores with only the formulation Eq. 9 is not ideal because the action of the inverse in Eq. 3 shall bring the same numerical issues as before into the problem. To overcome this, we derive a recursive formula for $\tilde{T}_{z}=\tilde{C}_{z}\tilde{U}_{z}$, which can be generalized to dimension-wise TT cores in the tensor setting. The recursive formula for $T$ reads
(10) $\tilde{T}_{z+1}=[\tilde{T}_{z}-\delta_{z}s_{z}\tilde{T}_{z}(i,:),-\delta_{z}s_{z}],$
@@
where $s_{z}=\tilde{T}_{z}A_{\mathcal{I},j}-A_{:,j}$ is a vector, and the derivation is immediate from Eq. 6. In terms of submatrix $(\mathcal{K},\mathcal{L})$, $\mathcal{L}$ contributes to the selection of rows $\tilde{R}_{z}$, so $\tilde{T}_{z}$ can be partitioned with respect to only $\mathcal{K}$:
(11) $\tilde{T}_{z+1}(\mathcal{K},:)=[\tilde{T}_{z}(\mathcal{K},:)-\delta_{z}s_{z}(\mathcal{K})\tilde{T}_{z}(i,:),-\delta_{z}s_{z}(\mathcal{K})],$
@@
where $s_{z}(\mathcal{K})=\tilde{T}_{z}(\mathcal{K},:)A_{\mathcal{I},j}-A_{\mathcal{K},j}$. This suggests that, instead of using the same 2D grid as in the pivot selection stage as depicted in Fig. 3, a 1D process grid is sufficient for the construction of $\tilde{T}_{z}$, so it is most convenient to use Eq. 11 once all pivots for the cross approximation are discovered. Algorithm 3 describes this procedure of building $T$ on a 1D process grid with all pivots selected, and we shall use this algorithm in the TT case for all dimensions. Furthermore, as the 2D process grid consists of much more processes than the 1D grid, we only need to select a few to build the 1D grid, as opposed to introducing more processes into our algorithm. We can also choose the processes so that communication to gain the information for building $T$ is minimal.
@@
Fig. 4 shows a simple illustration of applying Algorithm 3 to construct the approximation Eq. 11 on the same example in Fig. 3. Here, we also use a 2D grid of size $2\times 3$. Assume that we find a total of three pivots. The pivots are labeled as red crosses and are found on three processes. Then, as the row indices are partitioned to two pieces, we can use two processes $P_{1}$ and $P_{2}$ to construct $T(\mathcal{K},:)$. As the input of Algorithm 3, it is required that each process needs to know $A(\mathcal{K}\cup\mathcal{I},\mathcal{J})$. As depicted in the top middle plot in Fig. 4, $P_{1}$ needs to obtain blue lines and dots of $A$ from the 2D process grid; similarly, $P_{2}$ needs to obtain orange lines and dots. To minimize communication, we choose to use $P_{1,3}$ as $P_{1}$ and $P_{2,3}$ as $P_{2}$ since the column indices of two pivots are in the same column partitioning of $A$. The bottom two pictures show an example of the communications in lines 4-8 in Algorithm 3 that we handle the pivots one by one. Suppose we finish with two pivots and there is only one remaining (see bottom left), we first determine that the row index is in the possesion of $P_{1}$. This indicates that $P_{2}$ needs to receive the orange row vector from $P_{1}$ (see bottom right) to build its portion of $T$.
@@
Figure 4: The communication pattern of using Algorithm 3 to build approximation Eq. 11. Top left: three pivots are found with a $2\times 3$ grid and are denoted by red crosses. Top middle: matrix information (both lines and dots of two colors) needed for the construction of $T$. Top right: A 1D process grid used for building $T$ if row indices of $A$ are partitioned to two pieces. $P_{1}$ deals with the top half so it needs all blue entries from top middle plot. $P_{2}$ needs all orange elements from top middle plot. Process $P_{1,3}$ and $P_{2,3}$ from the 2D grid in top left plot are used as $P_{1}$ and $P_{2}$ respectively to minimize communication of getting the necessary elements. Bottom left: suppose Algorithm 3 is performed and there is only one pivot left. Bottom right: at this step, since $P_{1}$ holds the row index of the pivot, to build the orange portion of $T$, $P_{2}$ needs to receive the orange row from $P_{1}$.
@@
12:     Construct $T(\mathcal{K},:)=[T(\mathcal{K},:)-\delta s(\mathcal{K})T(i_{z},:)\quad-\!\delta s(\mathcal{K})]$.
@@
Below we analyze the communication complexity for Algorithm 3. Suppose we select $N$ pivots, then with some explicit counting, we can find that a process in the 1D grid needs to receive at most one $\delta$ value and a row vector of length $z$ in the $z$th iteration. Therefore, for all $N-1$ iterations, one process receives at most
(12) $(N-1)+\sum_{z=1}^{N-1}z=(N-1)(N/2+1)$
@@ 3.2.1 Pivot selection
Since cross approximation for the TT format requires pivots regarding multiple dimensions, we aim to discover pivots with the nestedness property [dolgov2020parallel], so that the approximation generated by Algorithm 1 recovers the exact same elements of the original tensor on positions $(\mathcal{I}_{\leq k},\mathbb{I}_{k+1},\mathcal{J}_{>k+1})$ throughout all iterations. Mathematically, the nestedness of the pivots can be represented as
@@
for $11\leq k\leq d-$ 1. In subtensor parallelism, we hope to maintain this nestedness so that running Algorithm 1 preserves interpolation on the entire tensor.
@@
Using the similar argument, we can show $\mathcal{L}_{>k}^{(c)}\subset\mathcal{L}_{>k+1}^{(c)}\times\mathcal{L}_{k+1}^{(c)}$ as well. These provide nestedness guarantee of the subtensor parallel algorithm. As a major corollary, the global best pivots can be thus obtained as the best of the local pivots on each subtensor. In practice, one could perform the pivot selection in a dimensional parallel manner. After pivots are selected for all $d-1$ tensor unfoldings, we are ready to carry out our next task, i.e. to build the TT cores Eq. 9 from the chosen pivots.
@@ 3.2.2 TT core construction
Figure 5: The general procedure of constructing an internal core at a global level.
The formation of the internal cores $\mathcal{G}_{k}$ for $12\leq k\leq d-$ 1 can start to suffer from computational expense, if we first form the full $T_{\leq k}$ and then extract the correct rows determined by the pivots in $\mathcal{I}_{\leq k-1}$. Instead, we implement the reverse by only computing what would be the extracted rows of $T_{\leq k}$. In other words, we effectively slice prior to the construction of $T_{\leq k}$ so that we only compute a submatrix of $T_{\leq k}$. This can be phrased as extraction of a submatrix which contributes to $T_{\leq k}$ from the full unfolding $X_{k}$. Since elements of $\mathcal{I}_{\leq k-1}$ populate the first $k-1$ indices of $\mathcal{X}$, we can make this extraction by computing $\mathcal{X}(\mathcal{I}_{\leq k-1},:)$. In practice, this can be done efficiently with the function np.ix_ in numpy by calling $\mathcal{X}[\text{np.ix}\_(\mathcal{I}_{\leq k-1})]$. This reduces the size of the matrix used for computation of $T_{\leq k}$ from size $n^{k}\times r$ to size $rn\times r$. In this setting, the neighboring indices $\mathcal{I}_{\leq k-1}$ must be communicated to the processes in charge of the computation of $\mathcal{G}_{k}$ before the start of construction. Once all processes that require index information for slicing have the necessary information, then we can use Algorithm 3 to compute the necessary submatrix of $T_{\leq k}$, which can then be gathered and reshaped into the global TT cores $\mathcal{G}_{k}$ using the following relation
(14) $\mathcal{G}_{k}=\begin{cases}\text{reshape}(T_{\leq 1},[1,n,r])&k=1\\ \text{reshape}(T_{\leq k},[r,n,r])&2\leq k\leq d-1\\ \text{reshape}(X_{d-1}(\mathcal{I}_{d-1},:),[r,n,1])&k=d\end{cases}.$
@@ 3.2.3 Grid development and overall algorithm
Then, we follow the discussion in Section 3.2.2 and use Algorithm 3 to build the TT cores. In particular, for dimension $k$, $T_{\leq k}$ from Section 3.2.2 is associated with the $k$th unfolding of the tensor, whose row indices represent the combined indices of the first $k$ dimensions of the tensor. Therefore, if each dimension is partitioned to $M$ pieces by the subtensor grid, then we need to build a 1D grid with $M^{k}$ partitions for $T_{\leq k}$. This 1D grid can be simply reshaped and understood as a $k$-dimensional grid (with $M$ pieces per dimension) as a higher-dimensional analogue of the 1D grid used in the matrix case in Section 3.1.3. In practice, we require that one process only appears at most once in any $k$-dimensional grid for $11\leq k\leq d-$ 1 for load balance, and the processes are chosen so that communications to switch between the grids for Algorithm 2 and Algorithm 3 are minimal.
@@
[…]
@@
Finally, we describe this subtensor TT cross approximation with iterative formulations in Algorithm 4.
@@
2:A TT approximation for $\mathcal{X}$ in distributed memory. Process $c$ returns portion of the TT core $\mathcal{G}_{k}^{(c)}$ if $c$ is used for constructing the $k$th TT core.
3:for $10\leq k\leq d-$ 1 do
@@
9:              Receive full $\mathcal{I}_{\leq k-1}$ from the root of the grid for dimension $k-1$.

### v2

@@ 3.1 Submatrix parallel matrix cross approximation
3.1.1 A derivation of an iterative construction of matrix cross approximation
In the following section, we go through the derivation of an iterative construction of matrix cross approximations. The final formula given in Eq. 8 is originally proposed in [6, Section 2], but we show a full proof for completeness. In addition, Eq. 11 in Section 3.1.3 is a corollary of the derivation, and to the best of our knowledge, it is a new formula for core construction in the tensor setting.
@@ 3.1.1 A derivation of an iterative construction of matrix cross approximation
where $\tilde{E}_{z}:=\tilde{A}_{z}-A$ is the difference between the target matrix $A$ and current approximation $\tilde{A}_{z}$. A similar form of Eq. 8 can be found in [13], where the construction is led by an LU decomposition of the approximation. In other words, the construction starts with full data of the original matrix. On the contrary, ours begins with a zero initialization of the approximation.
As an example, we perform a greedy cross approximation on a $\times100$ Hilbert matrix defined elementwise via
@@
which is shown to have rapidly decaying singular values in [7]. In Fig. 2, we compare the behaviors of Eq. 8 and the direct computation for approximation construction, where the evaluation of the inverse of the cross sections $A_{\mathcal{I},\mathcal{J}}$ is used. There we can see that the use of formula Eq. 8 provides better numerical stability up to machine precision, while the direct method suffers from degeneracy when selected rank is larger than 12. A separate advantage of Eq. 8 is that we are able to make global approximation updates at a local level, which is essential for communication efficiency of submatrix and subtensor algorithms discussed in the following subsections.
Figure 2: Relative Error versus rank using formula (8) and a direct computation of the factor matrices. The test was run a $\times100$ Hilbert matrix.
@@ 3.1.2 Iterative construction of submatrix cross approximation
2:A new pivot $(i^{*},j^{*})$ for the entire matrix AA.
@@ 3.1.3 An alternative iterative construction of matrix and submatrix cross approximation
For submatrix ACA, Eq. 10 is sufficient for both finding pivots and constructing approximations. However, building TT cores with only the formulation Eq. 10 is not ideal because the action of the inverse in Eq. 4 shall bring the same numerical issues as before into the problem. To overcome this, we derive a recursive formula for $\tilde{T}_{z}=\tilde{C}_{z}\tilde{U}_{z}$, which can be generalized to dimension-wise TT cores in the tensor setting as they take the form of $CU$. We start by taking the first term in Eq. 6 and expand and substitute $\tilde{T}_{z}$ where it appears.
$\tilde{C}_{z+1}\tilde{U}_{z+1}$ $=\begin{bmatrix}\tilde{C}_{z}(\tilde{U}_{z}+\delta_{z}\tilde{U}_{z}A_{\mathcal{I},j}A_{i,\mathcal{J}}\tilde{U}_{z})-\delta_{z}A_{:,j}A_{i,\mathcal{J}}\tilde{U}_{z},&-\delta_{z}\tilde{C}_{z}\tilde{U}_{z}A_{\mathcal{I},j}+\delta_{z}A_{:,j}\end{bmatrix}$
$=\begin{bmatrix}\tilde{T}_{z}+(\delta_{z}\tilde{T}_{z}A_{\mathcal{I},j}-\delta_{z}A_{:,j})A_{i,\mathcal{J}}\tilde{U}_{z},&-\delta_{z}\tilde{T}_{z}A_{\mathcal{I},j}+\delta_{z}A_{:,j}\end{bmatrix}$.
If we then define $s_{z}=\tilde{T}_{z}A_{\mathcal{I},j}-A_{:,j}$, and note that $A_{i,\mathcal{J}}\tilde{U}_{z}=\tilde{C}_{z}(i,:)\tilde{U}_{z}=\tilde{T}_{z}(i,:)$, then we have the recursive formula for $\tilde{T}_{z+1}$:
(11) $\tilde{T}_{z+1}=\left[\tilde{T}_{z}+\delta_{z}s_{z}\tilde{T}_{z}(i,:),\quad-\delta_{z}s_{z}\right].$
We include Fig. 4 to demonstrate the functionality of Eq. 11. In this figure, we see that at each iteration, the red block grows one column at a time, and it depends on the previous iterate via the formula $\tilde{T}_{z}+\delta_{z}s_{z}\tilde{T}_{z}(i,:)$. This new altered block is then concatenated on the right by a new column vector $\delta_{z}s_{z}$ (shown in blue), and then the full block matrix is used to complete the next step of iteration.
$\tilde{T}_{1}$$\tilde{T}_{2}$$\tilde{T}_{3}$$\tilde{T}_{k}$$\tilde{T}_{k}+\delta_{k}s_{k}\tilde{T}_{k}(i,:)$$\delta_{k}s_{k}$$\tilde{T}_{z}$
Figure 4: Visual representation of the use of Eq. 11 to construct $\tilde{T}_{z}$. At each iteration a new matrix is constructed by computing the red block using the first term of Eq. 11, and then the blue column given by the second term of Eq. 11 is concatenated on the right.
@@
In terms of submatrix $(\mathcal{K},\mathcal{L})$, $\mathcal{L}$ contributes to the selection of rows $\tilde{R}_{z}$, so $\tilde{T}_{z}$ can be partitioned with respect to only $\mathcal{K}$:
(12) $\tilde{T}_{z+1}(\mathcal{K},:)=\left[\tilde{T}_{z}(\mathcal{K},:)+\delta_{z}s_{z}(\mathcal{K})\tilde{T}_{z}(i,:),\quad-\delta_{z}s_{z}(\mathcal{K})\right],$
@@
where $s_{z}(\mathcal{K})=\tilde{T}_{z}(\mathcal{K},:)A_{\mathcal{I},j}-A_{\mathcal{K},j}$. We would like to emphasize that in Eq. 12, we only need one group of index sets $\mathcal{K}$ to partition the matrices $\tilde{T}_{z}$ into submatrices that contain all the columns. This suggests that, instead of using the same 2D grid as in the pivot selection stage as depicted in Fig. 3, a 1D process grid is sufficient for the construction of $\tilde{T}_{z}$, so it is most convenient to use Eq. 12 once all pivots for the cross approximation are discovered. Algorithm 4 describes this procedure of building TT on a 1D process grid with all pivots selected, and we shall use this algorithm in the TT case for all dimensions. Simply speaking, Algorithm 4 iteratively uses Eq. 12 to build a submatrix $T(\mathcal{K},:)$ of the final approximation $T$ in $A=TR$ on a distributed memory computing environment, via plugging the discovered pivots one at a time. Based on whether $i_{z}$, the row index of the used pivot, is in $\mathcal{K}$, this process either computes $\delta$ or receives $\delta$ and $T(i_{z},:)$ from another process. Furthermore, as the 2D process grid consists of much more processes than the 1D grid, we only need to select a few to build the 1D grid, as opposed to introducing more processes into our algorithm. We can also choose the processes so that communication to gain the information for building TT is minimal.
@@
Fig. 5 shows a simple illustration of applying Algorithm 4 to construct the approximation Eq. 12 on the same example in Fig. 3. Here, we also use a 2D grid of size $2\times 3$. Assume that we find a total of three pivots. The pivots are labeled as red crosses and are found on three processes. Then, as the row indices are partitioned to two pieces, we can use two processes $P_{1}$ and $P_{2}$ to construct $T(\mathcal{K},:)$. As the input of Algorithm 4, it is required that each process needs to know $A(\mathcal{K}\cup\mathcal{I},\mathcal{J})$. As depicted in the top middle plot in Fig. 5, $P_{1}$ needs to obtain blue lines and dots of $A$ from the 2D process grid; similarly, $P_{2}$ needs to obtain orange lines and dots. To minimize communication, we choose to use $P_{1,3}$ as $P_{1}$ and $P_{2,3}$ as $P_{2}$ since the column indices of two pivots are in the same column partitioning of $A$. The bottom two pictures show an example of the communications in lines 4-8 in Algorithm 4 that we handle the pivots one by one. Suppose we finish with two pivots and there is only one remaining (see bottom left), we first determine that the row index is in the possession of $P_{1}$. This indicates that $P_{2}$ needs to receive the orange row vector from $P_{1}$ (see bottom right) to build its portion of $T$.
@@
Figure 5: The communication pattern of using Algorithm 4 to build approximation Eq. 12. Top left: three pivots are found with a $2\times 3$ grid and are denoted by red crosses. Top middle: matrix information (both lines and dots of two colors) needed for the construction of $T$. Top right: A 1D process grid used for building TT if row indices of $A$ are partitioned to two pieces. $P_{1}$ deals with the top half so it needs all blue entries from top middle plot. $P_{2}$ needs all orange elements from top middle plot. Process $P_{1,3}$ and $P_{2,3}$ from the 2D grid in top left plot are used as $P_{1}$ and $P_{2}$ respectively to minimize communication of getting the necessary elements. Bottom left: suppose Algorithm 4 is performed and there is only one pivot left. Bottom right: at this step, since $P_{1}$ holds the row index of the pivot, to build the orange portion of $T$, $P_{2}$ needs to receive the orange row from $P_{1}$.
@@
12:  Construct $T(\mathcal{K},:)=[T(\mathcal{K},:)+\delta s(\mathcal{K})T(i_{z},:)\quad-\!\delta s(\mathcal{K})]$.
@@
Below we analyze the communication complexity for Algorithm 4. Suppose we select $N$ pivots, then with some explicit counting, we can find that a process in the 1D grid needs to receive at most one $\delta$ value and a row vector of length z1z-1 in the $z$th iteration. Therefore, for all N1N-1 iterations, one process receives at most
(13) $(N-1)+{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{\sum_{z=2}^{N}(z-1)}}=(N-1)(N/2+1)$
@@ 3.2.1 Pivot selection
Since cross approximation for the TT format requires pivots regarding multiple dimensions, we aim to discover pivots with the nestedness property [17], so that the approximation generated by Algorithm 2 recovers the exact same elements of the original tensor on positions $(\mathcal{I}_{\leq k},\mathbb{I}_{k+1},\mathcal{J}_{>k+1})$ throughout all iterations. For a more detailed discussion of how nestedness preserves interpolation and why the greedy pivot selection strategy attains nestedness, we refer the readers to [17, 41]. Mathematically, the nestedness of the pivots can be represented as
@@
for $1\leq k\leq d-1$. In subtensor parallelism, we also hope to maintain this nestedness by using Algorithm 2 on subtensors instead of the entire tensor.
@@
Using the similar argument, we can show $\mathcal{L}_{>k}^{(c)}\subset\mathcal{L}_{>k+1}^{(c)}\times\mathcal{L}_{k+1}^{(c)}$ as well. These provide nestedness guarantee of the subtensor parallel algorithm. As a major corollary, the global best pivots can be thus obtained as the best of the local pivots on each subtensor. In practice, one could perform the pivot selection in a dimensional parallel manner. After pivots are selected for all d1d-1 tensor unfoldings, we are ready to carry out our next task, i.e. to build the TT cores Eq. 10 from the chosen pivots.
@@ 3.2.2 TT core construction
Figure 6: The general procedure of constructing an internal core at a global level. The first image depicts the full superblock $X_{k}(\mathcal{I}_{\leq k-1}\otimes\mathbb{I}_{k},\mathbb{I}_{k+1}\otimes\mathcal{J}_{>k+1})$ for $X_{k}$ in the blue submatrices. In this image we also show the specific columns $\mathcal{J}_{>k}$ in red, which are sliced out of the superblock to get the second image depicting $X_{k}(\mathcal{I}_{\leq k-1}\otimes\mathbb{I}_{k},\mathcal{J}_{>k})$. This data is given to Eq. 11 (see Fig. 4) to obtain $T_{\leq k}$. Finally, $T_{\leq k}$ is reshaped to $\mathcal{G}_{k}$.
The formation of the internal cores $\mathcal{G}_{k}$ for $2\leq k\leq d-1$ can start to suffer from computational expense, if we first form the full $T_{\leq k}$ and then extract the correct rows determined by the pivots in $\mathcal{I}_{\leq k-1}$. Instead, we implement the reverse by only computing what would be the extracted rows of $T_{\leq k}$. In other words, we effectively slice prior to the construction of $T_{\leq k}$ so that we only compute a submatrix of $T_{\leq k}$. This can be phrased as extraction of a submatrix which contributes to $T_{\leq k}$ from the full unfolding $X_{k}$. Since elements of $\mathcal{I}_{\leq k-1}$ populate the first k1k-1 indices of $\mathcal{X}$, we can make this extraction by computing $\mathcal{X}(\mathcal{I}_{\leq k-1},:)$. In practice, this can be done efficiently with the function np.ix_ in numpy by calling $\mathcal{X}[\text{np.ix}\_(\mathcal{I}_{\leq k-1})]$. This reduces the size of the matrix used for computation of $T_{\leq k}$ from size $n^{k}\times r$ to size $rn\times r$. In this setting, the neighboring indices $\mathcal{I}_{\leq k-1}$ must be communicated to the processes in charge of the computation of $\mathcal{G}_{k}$ before the start of construction. Once all processes that require index information for slicing have the necessary information, then we can use Algorithm 4 to compute the necessary submatrix of $T_{\leq k}$, which can then be gathered and reshaped into the global TT cores $\mathcal{G}_{k}$ using the following relation
(15) $\mathcal{G}_{k}=\begin{cases}\text{reshape}(T_{\leq 1},[1,n,r])&k=1\\ \text{reshape}(T_{\leq k},[r,n,r])&2\leq k\leq d-1\\ \text{reshape}(\mathcal{X}(\mathcal{I}_{d-1},:),[r,n,1])&k=d\end{cases}.$
@@ 3.2.3 Grid development and overall algorithm
Then, we follow the discussion in Section 3.2.2 and use Algorithm 4 to build the TT cores. In particular, for dimension $k$, $T_{\leq k}$ from Section 3.2.2 is associated with the $k$th unfolding of the tensor, whose row indices represent the combined indices of the first $k$ dimensions of the tensor. Therefore, if each dimension is partitioned to $M$ pieces by the subtensor grid, then we need to build a 1D grid with $M^{k}$ partitions for $T_{\leq k}$. This 1D grid can be simply reshaped and understood as a $k$-dimensional grid (with $M$ pieces per dimension) as a higher-dimensional analogue of the 1D grid used in the matrix case in Section 3.1.3. In practice, we require that one process only appears at most once in any $k$-dimensional grid for $1\leq k\leq d-1$ for load balance, and the processes are chosen so that communications to switch between the grids for Algorithm 3 and Algorithm 4 are minimal.
@@
$\mathcal{X}$$\mathcal{K}_{1}$$\mathcal{K}_{2}$$\mathcal{K}_{3}$$P_{1}$$P_{2}$$P_{3}$$P_{4}$$P_{5}$$P_{7}$$P_{8}$$T_{\leq 1}$$\mathcal{K}_{1}$$P_{2}$$P_{1}$$T_{\leq 2}$$\mathcal{K}_{1}$$\mathcal{K}_{2}$$P_{3}$$P_{4}$$P_{5}$$P_{6}$
@@
Finally, we describe this subtensor TT cross approximation with iterative formulations in Algorithm 5. In practice, Algorithm 5 is performed with multiple process grids, and more details about data transfers and communication patterns can be found in the next section.
@@
2:A TT approximation for $\mathcal{X}$ in distributed memory. Process $c$ returns portion of the TT core $\mathcal{G}_{k}^{(c)}$ if $c$ is used for constructing the $k$th $T$ core.
3:for $0\leq k\leq d-1$ do
@@
9:     Receive full $\mathcal{I}_{\leq k-1}$ from the root of the grid for dimension k1k-1.